# LtU-ILI: An All-in-One Framework for Implicit Inference in Astrophysics and Cosmology

### v1

@@ Abstract
This paper presents the Learning the Universe Implicit Likelihood Inference (LtU-ILI) pipeline, a codebase for rapid, user-friendly, and cutting-edge machine learning (ML) inference in astrophysics and cosmology. The pipeline includes software for implementing various neural architectures, training schema, priors, and density estimators in a manner easily adaptable to any research workflow. It includes comprehensive validation metrics to assess posterior estimate coverage, enhancing the reliability of inferred results. Additionally, the pipeline is easily parallelizable, designed for efficient exploration of modeling hyperparameters. To demonstrate its capabilities, we present real applications across a range of astrophysics and cosmology problems, such as: estimating galaxy cluster masses from X-ray photometry; inferring cosmology from matter power spectra and halo point clouds; characterising progenitors in gravitational wave signals; capturing physical dust parameters from galaxy colors and luminosities; and establishing properties of semi-analytic models of galaxy formation. We also include exhaustive benchmarking and comparisons of all implemented methods as well as discussions about the challenges and pitfalls of ML inference in astronomical sciences. All code and examples are made publicly available at https://github.com/maho3/ltu-ili.
@@ 1 Introduction
Statistical inference of unknown quantities is a fundamental problem in science. In the astronomical sciences, we largely rely on Bayesian inference to test new physical models (Feigelson and Babu(2012); Dodelson and Schmidt(2020); Eadie et al.(2023)Eadie, Speagle, Cisewski-Kehe, Foreman-Mackey, Huppenkothen, Jones, Springford, and Tak), wherein we assume some prior hypothesis of an observed system and calculate constraints on model parameters which would be consistent with our observations. For nearly a century, the practice of building linear or perturbative physical models from first-principles allowed us to make substantial progress towards understanding the universe. Yet, as highlighted in the 2020 Decadal Survey (National Academies of Sciences, Engineering, and Medicine(2023)), exploring the complex, nonlinear regime of physical phenomena through data-driven inference promises significant gains in constraining power. The large data volume of next-generation surveys, the improvements in high-resolution simulations, and the rapid rise of machine learning (ML) techniques have driven an explosion of inquiry into how one can automatically and rapidly learn complex physical phenomena (Carleo et al.(2019)Carleo, Cirac, Cranmer, Daudet, Schuld, Tishby, Vogt-Maranto, and Zdeborová). Despite the growing enthusiasm, a major hurdle to the widespread adoption of data science methods in astronomy is in their limited accessibility to the general community, which when compounded by the rapid pace of developments in the field, makes it challenging for individuals to stay well-informed on best practices (Huppenkothen et al.(2023)Huppenkothen, Ntampaka, Ho, Fouesneau, Nord, Peek, Walmsley, Wu, Avestruz, Buck et al.). There is currently no widespread consensus within the field regarding the approach to building a framework for implicit inference problems that is both reliable, robust, and accessible.
@@
Implicit Likelihood Inference (ILI, Cranmer et al.(2020)Cranmer, Brehmer, and Louppe; Marin et al.(2012)Marin, Pudlo, Robert, and Ryder), also known as simulation-based inference (SBI) and likelihood-free inference (LFI), is an approach for learning the statistical relationship between parameters and data. ILI is a mathematically rigorous way of framing ML under the umbrella of Bayesian statistics, a language common to astrophysicists and cosmologists since the turn of the century (Christensen et al.(2001)Christensen, Meyer, Knox, and Luey). It aims to estimate a full posterior distribution over parameters, including predictive uncertainties. It is best explained in contrast to traditional explicit likelihood analyses, wherein one might write down an analytic likelihood to represent the probability of observations given a model and its parameters, and then use sampling methods such as Markov Chain Monte Carlo (MCMC; Brooks et al.(2011)Brooks, Gelman, Jones, and Meng) to sample from the posterior. In ILI, one seeks to automatically learn the likelihood, i.e. to model the relationship between model parameters and observations through training on simulations. As such, ILI can be applied to infer parameters for any phenomenon that can be simulated, without the need for assumptions about the analytical form of the likelihood. Furthermore, ILI can accommodate complex data cuts (provided that consistent filters are applied to both the data and the forward model) something that can be hard to model in likelihood-based approaches.
@@
Given a prescribed training set of model parameters and observations or a fast physical simulator, ILI can produce robust, tightly-constraining inference without analytic likelihoods. In astronomy and cosmology, ILI has consistently been shown to accelerate and improve upon traditional analyses such as cosmological galaxy lensing and clustering (e.g., Jeffrey et al.(2021)Jeffrey, Alsing, and Lanusse; Makinen et al.(2021)Makinen, Charnock, Alsing, and Wandelt; Makinen et al.(2022a)Makinen, Charnock, Lemos, Porqueres, Heavens, and Wandelt; de Santi et al.(2023a)de Santi, Shao, Villaescusa-Navarro, Abramo, Teyssier, Villanueva-Domingo, Ni, Anglés-Alcázar, Genel, Hernandez-Martinez et al.; Hahn et al.(2023)Hahn, Lemos, Parker, Blancard, Eickenberg, Ho, Hou, Massara, Modi, Dizgah et al.), gravitational waves (e.g., Dax et al.(2021a)Dax, Green, Gair, Macke, Buonanno, and Schölkopf; Cheung et al.(2022)Cheung, Wong, Hannuksela, Li, and Ho), galaxy cluster mass estimation (e.g., Ho et al.(2022)Ho, Ntampaka, Rau, Chen, Lansberry, Ruehle, and Trac; de Andres et al.(2022)de Andres, Cui, Ruppin, De Petris, Yepes, Gianfagna, Lahouli, Aversano, Dupuis, Jarraya et al.), galaxy morphology (e.g., Walmsley et al.(2020)Walmsley, Smith, Lintott, Gal, Bamford, Dickinson, Fortson, Kruk, Masters, Scarlata et al.; Ghosh et al.(2022)Ghosh, Urry, Rau, Perreault-Levasseur, Cranmer, Schawinski, Stark, Tian, Ofman, Ananna et al.), stellar streams (e.g., Hermans et al.(2021)Hermans, Banik, Weniger, Bertone, and Louppe; Alvey et al.(2023)Alvey, Gerdes, and Weniger), and exoplanets (e.g., Rogers et al.(2023)Rogers, Janó Muñoz, Owen, and Makinen; Aubin et al.(2023)Aubin, Cuesta-Lazaro, Tregidga, Viaña, Garraffo, Gordon, López-Morales, Hargreaves, Makhnev, Drake et al.).
@@
Despite its popularity, robust applications of ILI have their challenges. First, fully Bayesian models include capturing epistemic (modeling) uncertainty, which requires marginalizing over model parameter uncertainty (i.e., network weights and biases) and is intractable for large architectures. However, approximations to this method using dropout marginalization (Gal and Ghahramani(2016)), model ensembling (Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell) or stochastic weight averaging (Maddox et al.(2019)Maddox, Izmailov, Garipov, Vetrov, and Wilson; Lemos et al.(2023)Lemos, Cranmer, Abidi, Hahn, Eickenberg, Massara, Yallup, and Ho) prove to be empirically sufficient surrogates. Second, choices of modeling hyperparameters (e.g., preprocessing, architecture, learning procedures, etc.) can greatly affect the quality of a trained ILI model. This is generally addressed through model ensembling or hyperparameter searches (e.g., Jin et al.(2019)Jin, Song, and Hu; White et al.(2021)White, Neiswanger, and Savani). Lastly, a pervasive problem in ILI is model misspecification, wherein the simulator or training catalog used to learn likelihoods is not representative of reality. Methods such as SBI++ (Wang et al.(2023)Wang, Leja, Villar, and Speagle) deal with missing data and outliers and present applications to galactic photometric redshift inference (Modi et al.(2023)Modi, Pandey, Ho, Hahn, Blancard, and Wandelt). For a thorough discussion of caveats, see Section 6. We argue that addressing these common problems is paramount to responsible application of ILI, and testing them exhaustively necessitates a unified, accessible framework.
@@
At the time of publishing, LtU-ILI offers the broadest and most complete toolbox of features among ILI software. Over sbi, we include the capacity for custom data-loading procedures, additional neural density estimators (NDEs), and adaptability to exotic embedding architectures (including sequential and graph encoders). Over pydelfi, we allow the capacity for neural posterior and ratio estimation. Over lampe, we provide a configuration-based interface to training procedures, the ability to specify separate proposal and prior distributions, and automated validation tests. We also include new multivariate coverage tests such as TARP (Lemos et al.(2023)Lemos, Coogan, Hezaveh, and Levasseur) and adaptive, automatic integration of data-loading, inference, and validation. We also remark that, during the development of this work, nbi (Zhang et al.(2023)Zhang, Bloom, and Hernitschek) was released as a neural posterior estimation toolkit focused on spectra and light-curve analysis. Though we do not natively integrate nbi, we remark that we encompass the functionality of nbi in our various neural posterior estimation implementations and extend them with alternatives and additional features.
In this software release paper, we provide background, details, and example applications of the LtU-ILI pipeline. The code is publicly available on Github [^1] https://github.com/maho3/ltu-ili. In Section 2, we review the theoretical foundations of ILI. In Section 3, we describe the structure of our pipeline and the various configurations that users can interact with. In Section 4, we apply LtU-ILI on several synthetic problems where the true solution is known, to demonstrate and benchmark its performance relative to traditional methods. In Section 5, we provide numerous diverse examples of parameter inference in science using LtU-ILI. In 6, we discuss the important considerations to take into account when designing an ILI pipeline, as well as possible failure modes and relevant solutions. Lastly, we provide a summary of our findings in Section 7.
@@ 2 Implicit Likelihood Inference
There are two approaches to modeling the terms in Equation 1: explicit (theory-driven) analyses or implicit (data-driven) analyses. In a classical explicit likelihood-based analysis, we assume an analytic form for the likelihood, $\mathcal{P}\left(\bm{x}|\bm{\theta}\right)$, and substitute it into Equation 1. Typically, this expression is derived from analytic theory or simple simulations, coupled with strong assumptions on the statistical distribution of noise or unknown physics. Sometimes, these statistical assumptions are relatively simple, e.g., a Gaussian distribution for continuous data (e.g., Tegmark et al.(2004)Tegmark, Blanton, Strauss, Hoyle, Schlegel, Scoccimarro, Vogeley, Weinberg, Zehavi, Berlind, Budavari, Connolly, Eisenstein, Finkbeiner, Frieman, Gunn, Hamilton, Hui, Jain, Johnston, Kent, Lin, Nakajima, Nichol, Ostriker, Pope, Scranton, Seljak, Sheth, Stebbins, Szalay, Szapudi, Verde, Xu, Annis, Bahcall, Brinkmann, Burles, Castander, Csabai, Loveday, Doi, Fukugita, Gott, Hennessy, Hogg, Ivezić, Knapp, Lamb, Lee, Lupton, McKay, Kunszt, Munn, O’Connell, Peoples, Pier, Richmond, Rockosi, Schneider, Stoughton, Tucker, Vanden Berk, Yanny, York, and SDSS Collaboration; Christensen and Meyer(2022); Porqueres et al.(2022)Porqueres, Heavens, Mortlock, and Lavaux) or a Poisson distribution for discrete counts (e.g., Boese and Doebereiner(2001); Braun et al.(2008)Braun, Dumm, De Palma, Finley, Karle, and Montaruli; Tsaprazi et al.(2022)Tsaprazi, Nguyen, Jasche, Schmidt, and Lavaux). Although these may be appropriate distributions for many situations, for a general inference problem we may not know which distribution to choose or whether this choice is valid. Also, MCMC-like samplers scale poorly with dimensionality and have bad mode coverage. This can cause problems for high-dimensional inference problems, especially if the likelihood is not differentiable with respect to the parameters. However, if we have access to a set of examples of the data-parameter relationship (i.e., samples from the joint distribution, $(\bm{x},\bm{\theta})\sim\mathcal{P}(\bm{x},\bm{\theta})$), we can use implicit likelihood analyses to learn the shape of a likelihood or posterior automatically with full differentiability. This flexibility makes ILI an extremely powerful tool for modeling complex or otherwise poorly-understood physical phenomena.
@@
ILI learns approximate posteriors, $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$, or likelihoods from the distribution of example data-parameter pairs in a training catalog, $\mathcal{D}_{\rm train}:=\{(\bm{x}_{i},\bm{\theta}_{i})\}_{i=1}^{N_{\rm train}}$. These data can take any form; they can be sensor readings in an experiment (e.g., Dingeldein et al.(2023)Dingeldein, Cossio, and Covino), images of the night sky (e.g., Ntampaka et al.(2019)Ntampaka, ZuHone, Eisenstein, Nagai, Vikhlinin, Hernquist, Marinacci, Nelson, Pakmor, Pillepich et al.), or even compressed summaries of high-dimensional stochastic properties (e.g., Modi et al.(2023)Modi, Pandey, Ho, Hahn, Blancard, and Wandelt). They can be gathered from a pre-run simulation (e.g., Fluri et al.(2022)Fluri, Kacprzak, Lucchi, Schneider, Refregier, and Hofmann), be simulated on-the-fly (e.g., Alsing et al.(2019)Alsing, Charnock, Feeney, and Wandelt), or curated from manually-labeled observations (e.g., Lanusse et al.(2018)Lanusse, Ma, Li, Collett, Li, Ravanbakhsh, Mandelbaum, and Póczos), but they must be representative of the true underlying problem. Given these data-parameter pairs, we wish to construct a model capable of doing inference on real data, i.e., to evaluate the posterior at some observed value $\bm{x}=\bm{x}_{\rm o}$.
@@ 2.1 Approximate Bayesian Computation
Approximate Bayesian Computation (ABC; Rubin(1984)) is the original basis of ILI. Given a prior $\mathcal{P}(\bm{\theta})$, a simulator ($\bm{\theta}\rightarrow\bm{x}$), and an observation $\bm{x}_{\rm o}$, ABC can construct a posterior. The general procedure of Rejection ABC is straightforward: (1) first, draw candidate parameter values from the prior $\bm{\theta}\sim\mathcal{P}(\bm{\theta})$, (2) forward-simulate these parameter values to observations $\bm{x}\sim\mathcal{P}(\bm{x}|\bm{\theta})$, (3) measure a distance metric between the candidate observations and the real observation, $d(\bm{x},\bm{x}_{\rm o})$, and (4) if the distance is sufficiently close to the real observation (i.e., $d(\bm{x},\bm{x}_{\rm o})<\epsilon$ for small $\epsilon$), accept the candidate parameters $\bm{\theta}$ as samples from the inferred posterior. ABC has been widely adopted in statistical inference because it does not require explicit knowledge of the likelihood and the accepted samples are guaranteed to converge to the true posterior as $\epsilon\rightarrow 0$. When $\epsilon$ is non-zero, the inferred posterior is guaranteed to be broader than the true posterior, enabling conservative inference. Despite successful applications in astrophysics and cosmology (e.g., Schafer and Freeman(2012); Cameron and Pettitt(2012); Hahn et al.(2017)Hahn, Vakili, Walsh, Hearin, Hogg, and Campbell), the ultimate problem with ABC methods is that the acceptance rate vanishes exponentially as the dimensionality of $\bm{\theta}$ increases, requiring thus significantly more simulations, which can be alleviated with a larger $\epsilon$ at the cost of targeting a broader posterior than the true posterior (Alsing et al.(2018)Alsing, Wandelt, and Feeney). This limitation makes ABC highly intractable in regimes where running simulations is expensive. Occasionally ABC is used as a synonym for LFI/SBI/ILI, however throughout the paper we exclusively use ABC to refer to the algorithm discussed above.

### v2

@@ Abstract
This paper presents the Learning the Universe Implicit Likelihood Inference (LtU-ILI) pipeline, a codebase for rapid, user-friendly, and cutting-edge machine learning (ML) inference in astrophysics and cosmology. The pipeline includes software for implementing various neural architectures, training schemata, priors, and density estimators in a manner easily adaptable to any research workflow. It includes comprehensive validation metrics to assess posterior estimate coverage, enhancing the reliability of inferred results. Additionally, the pipeline is easily parallelizable and is designed for efficient exploration of modeling hyperparameters. To demonstrate its capabilities, we present real applications across a range of astrophysics and cosmology problems, such as: estimating galaxy cluster masses from X-ray photometry; inferring cosmology from matter power spectra and halo point clouds; characterizing progenitors in gravitational wave signals; capturing physical dust parameters from galaxy colors and luminosities; and establishing properties of semi-analytic models of galaxy formation. We also include exhaustive benchmarking and comparisons of all implemented methods as well as discussions about the challenges and pitfalls of ML inference in astronomical sciences. All code and examples are made publicly available at https://github.com/maho3/ltu-ili.
@@ 1 Introduction
Statistical inference of unknown quantities is a fundamental problem in science. In the astronomical sciences, we largely rely on Bayesian inference to test new physical models (Feigelson and Babu, 2012; Dodelson and Schmidt, 2020; Eadie et al., 2023), wherein we assume some prior hypothesis of an observed system and calculate constraints on model parameters which would be consistent with our observations. For nearly a century, the practice of building linear or perturbative physical models from first-principles allowed us to make substantial progress towards understanding the universe. Yet, as highlighted in the 2020 Decadal Survey (National Academies of Sciences, Engineering, and Medicine, 2023), exploring the complex, nonlinear regime of physical phenomena through data-driven inference promises significant gains in constraining power. The large data volume of next-generation surveys, the improvements in high-resolution simulations, and the rapid rise of machine learning (ML) techniques have driven an explosion of inquiry into how one can automatically and rapidly learn complex physical phenomena (Carleo et al., 2019). Despite the growing enthusiasm, a major hurdle to the widespread adoption of data science methods in astronomy is in their limited accessibility to the general community, which when compounded by the rapid pace of developments in the field, makes it challenging for individuals to stay well-informed on best practices (Huppenkothen et al., 2023). There is currently no widespread consensus within the field regarding the approach to building a framework for implicit inference problems that is both reliable, robust, and accessible.
@@
Implicit Likelihood Inference (ILI, Cranmer et al., 2020; Marin et al., 2012), also known as simulation-based inference (SBI) and likelihood-free inference (LFI), is an approach for learning the statistical relationship between parameters and data. ILI is a mathematically rigorous way of framing ML under the umbrella of Bayesian statistics, a language common to astrophysicists and cosmologists since the turn of the century (Christensen et al., 2001). Given a parameterized model and observed data, ILI estimates posterior distributions over model parameters, while accounting for predictive uncertainties. It is best explained in contrast to traditional explicit likelihood analyses, wherein one might write down an analytic likelihood to represent the probability of observations given a model and its parameters, and then use sampling methods such as Markov Chain Monte Carlo (MCMC; Brooks et al., 2011) to sample from the posterior. In ILI, one seeks to automatically learn the likelihood, i.e. to model the relationship between model parameters and observations through training on simulations. As such, ILI can be applied to infer parameters for any phenomenon that can be simulated, without the need for assumptions about the analytical form of the likelihood. Furthermore, ILI can accommodate complex data cuts (provided that consistent filters are applied to both the data and the forward model) something that can be hard to model in likelihood-based approaches.
@@
Given a prescribed training set of model parameters and observations or a fast physical simulator, ILI can produce robust, tightly-constraining inference without analytic likelihoods. In astronomy and cosmology, ILI has consistently been shown to accelerate and improve upon traditional analyses such as cosmological galaxy lensing and clustering (e.g., Jeffrey et al., 2021; Makinen et al., 2021, 2022a; de Santi et al., 2023a; Hahn et al., 2023a), gravitational waves (e.g., Dax et al., 2021a; Cheung et al., 2022), galaxy cluster mass estimation (e.g., Ho et al., 2022; de Andres et al., 2022), galaxy morphology (e.g., Walmsley et al., 2020; Ghosh et al., 2022), stellar streams (e.g., Hermans et al., 2021; Alvey et al., 2023), and exoplanets (e.g., Rogers et al., 2023; Aubin et al., 2023).
@@
Despite its popularity, robust applications of ILI have their challenges. First, fully Bayesian models include capturing epistemic (modeling) uncertainty, which requires marginalizing over model parameter uncertainty (i.e., network weights and biases) and is intractable for large architectures. However, approximations to this method using dropout marginalization (Gal and Ghahramani, 2016), model ensembling (Lakshminarayanan et al., 2017) or stochastic weight averaging (Maddox et al., 2019; Lemos et al., 2023) prove to be empirically sufficient surrogates. Second, choices of modeling hyperparameters (e.g., preprocessing, architecture, learning procedures, etc.) can greatly affect the quality of a trained ILI model. This is generally addressed through model ensembling or hyperparameter searches (e.g., Jin et al., 2019; White et al., 2021). Lastly, a pervasive problem in ILI is model misspecification, wherein the simulator or training catalog used to learn likelihoods is not representative of reality. Methods such as SBI++ (Wang et al., 2023) deal with missing data and outliers and present applications to galactic photometric redshift inference (Modi et al., 2023). For a thorough discussion of caveats, see Section 6. We argue that addressing these common problems is paramount to responsible application of ILI, and testing them exhaustively necessitates a unified, accessible framework.
@@
  • The code includes state-of-the-art neural architectures, validation tools, and samplers for enabling ILI.
  • The pipeline is modular, easily customizable, and parallelizable for rapid testing of design choices and hyperparameters.
  • The interface is accessible to non-specialists while being practical for high-level production.
  • The methodologies automatically implement standard best practices for machine learning (See Huppenkothen et al., 2023, for a review).
@@
  • We unify several leading implicit inference codes–sbi (Tejero-Cantero et al., 2020), pydelfi (Alsing et al., 2019), and lampe (Rozet et al., 2021)–under a single common interface, for the first time allowing rigorous apples-to-apples comparisons of their performance.
  • We supply extensions to these codes, enabling them to utilize new types of embedding networks (e.g., for image- and graph-like datasets) and providing state-of-the-art validation metrics for enabling robust, maximally-informative inference.
  • We provide an accessible, comprehensive, dual Jupyter and command-line interface, for rapid research and development.
  • We demonstrate the abilities of ILI in an extensive selection of inference problems in astrophysics and cosmology, and we provide these as public benchmarks for future works.
At the time of publishing, LtU-ILI offers the broadest and most complete toolbox of features among ILI software. Over sbi, we include the capacity for custom data-loading procedures, additional neural density estimators (NDEs), and adaptability to exotic embedding architectures (including sequential and graph encoders). Over pydelfi, we allow the capacity for neural posterior and ratio estimation. Over lampe, we provide a configuration-based interface to training procedures, the ability to specify separate proposal and prior distributions, and automated validation tests. We also include new multivariate coverage tests such as TARP (Lemos et al., 2023) and adaptive, automatic integration of data-loading, inference, and validation.
In this software release paper, we provide background, details, and example applications of the LtU-ILI pipeline. The code is publicly available on Github [^1] https://github.com/maho3/ltu-ili. In Section 2, we review the theoretical foundations of ILI. In Section 3, we describe the structure of our pipeline and the various configurations that users can interact with. In Section 4, we apply LtU-ILI on several synthetic problems where the true solution is known, to demonstrate and benchmark its performance relative to traditional methods. In Section 5, we provide numerous diverse examples of parameter inference in science using LtU-ILI. In Section 6, we discuss the important considerations to take into account when designing an ILI pipeline, as well as possible failure modes and relevant solutions. Lastly, we provide a summary of our findings in Section 7.
@@ 2 Implicit Likelihood Inference
There are two approaches to modeling the terms in Equation 1: explicit (theory-driven) analyses or implicit (data-driven) analyses. In a classical explicit likelihood-based analysis, we assume an analytic form for the likelihood, $\mathcal{P}\left(\bm{x}|\bm{\theta}\right)$, and substitute it into Equation 1. Typically, this expression is derived from analytic theory or simple simulations, coupled with strong assumptions on the statistical distribution of noise or unknown physics. Sometimes, these statistical assumptions are relatively simple, e.g., a Gaussian distribution for continuous data (e.g., Tegmark et al., 2004; Christensen and Meyer, 2022; Porqueres et al., 2022) or a Poisson distribution for discrete counts (e.g., Boese and Doebereiner, 2001; Braun et al., 2008; Tsaprazi et al., 2023). Although these may be appropriate distributions for many situations, for a general inference problem we may not know which distribution to choose or whether this choice is valid. Also, MCMC-like samplers scale poorly with dimensionality and have bad mode coverage. This can cause problems for high-dimensional inference problems, especially if the likelihood is not differentiable with respect to the parameters. However, if we have access to a set of examples of the data-parameter relationship (i.e., samples from the joint distribution, $(\bm{x},\bm{\theta})\sim\mathcal{P}(\bm{x},\bm{\theta})$), we can use implicit likelihood analyses to learn the shape of a likelihood or posterior automatically with full differentiability. This flexibility makes ILI an extremely powerful tool for modeling complex or otherwise poorly-understood physical phenomena.
@@
ILI learns approximate posteriors $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$ or likelihoods $\hat{\mathcal{P}}(\bm{x}|\bm{\theta})$ from the distribution of example data-parameter pairs in a training catalog, $\mathcal{D}_{\rm train}:=\{(\bm{x}_{i},\bm{\theta}_{i})\}_{i=1}^{N_{\rm train}}$. These data can take any form; they can be sensor readings in an experiment (e.g., Dingeldein et al., 2023), images of the night sky (e.g., Ntampaka et al., 2019), or even compressed summaries of high-dimensional stochastic properties (e.g., Modi et al., 2023). They can be gathered from a pre-run simulation (e.g., Fluri et al., 2022), be simulated on-the-fly (e.g., Alsing et al., 2019), or curated from manually-labeled observations (e.g., Lanusse et al., 2018), but they must be representative of the true underlying problem. Given these data-parameter pairs, we wish to construct a model capable of doing inference on real data, i.e., to evaluate the posterior at some observed value $\bm{x}=\bm{x}_{\rm o}$.
@@ 2.1 Approximate Bayesian Computation
Approximate Bayesian Computation (ABC; Rubin, 1984) is the original basis of ILI. Given a prior $\mathcal{P}(\bm{\theta})$, a simulator ($\bm{\theta}\rightarrow\bm{x}$), and an observation $\bm{x}_{\rm o}$, ABC can construct a posterior. The general procedure of Rejection ABC is straightforward: (1) first, draw candidate parameter values from the prior $\bm{\theta}\sim\mathcal{P}(\bm{\theta})$, (2) forward-simulate these parameter values to observations $\bm{x}\sim\mathcal{P}(\bm{x}|\bm{\theta})$, (3) measure a distance metric between the candidate observations and the real observation, $d(\bm{x},\bm{x}_{\rm o})$, and (4) if the distance is sufficiently close to the real observation (i.e., $d(\bm{x},\bm{x}_{\rm o})<\epsilon$ for small $\epsilon$), accept the candidate parameters $\bm{\theta}$ as samples from the inferred posterior. ABC has been widely adopted in statistical inference because it does not require explicit knowledge of the likelihood and the accepted samples are guaranteed to converge to the true posterior as $\epsilon\rightarrow 0$. When $\epsilon$ is non-zero, the inferred posterior is guaranteed to be broader than the true posterior, enabling conservative inference. Despite successful applications in astrophysics and cosmology (e.g., Schafer and Freeman, 2012; Cameron and Pettitt, 2012; Hahn et al., 2017), the ultimate problem with ABC methods is that the acceptance rate vanishes exponentially as the dimensionality of $\bm{\theta}$ increases, requiring thus significantly more simulations, which can be alleviated with a larger $\epsilon$ at the cost of targeting a broader posterior than the true posterior (Alsing et al., 2018). This limitation makes ABC highly intractable in regimes where running simulations is expensive. Occasionally ABC is used as a synonym for LFI/SBI/ILI, however throughout the paper we exclusively use ABC to refer to the algorithm discussed above.